# Hypercube-RAG: Hypercube-Based Retrieval-Augmented Generation for In-domain Scientific Question-Answering

### v1

@@ Appendix F Ablation Study on Hypercube Dimensions
Appendix G Document Contents
| Document ID | Document Content |
| 565 | …Why is September the peak month for hurricanes? NASA oceanographer, Bill Patzert at the Jet Propulsion Laboratory, Pasadena, Calif., provided the answer: “Hurricanes are fueled by warm ocean temperatures and September is the end of the Northern Hemisphere ocean warming season. The 2008 Atlantic hurricane season started early with the formation of Tropical Storm Arthur on May 30, from the remnants of the eastern Pacific Ocean’s first storm, Alma, which crossed Central America and reformed in the Gulf of Mexico. It took one month and four days for the next storm to form, Bertha. Now, the action is definitely picking up. The tropical Atlantic is warm, but not unusually so.” Once a powerful Category 3 hurricane, now a tropical depression, Gustav moved from northwest Louisiana into northeastern Texas and into Arkansas by Sept. 3. Like Tropical Storm Fay in August, Gustav’s legacy will lie in large rainfall totals. According to the National Hurricane Center discussion on Sept. 2, "Storm total rainfalls are expected to be five to ten Inches with isolated maximums of 15 inches over portions of Louisiana, Arkansas and Mississippi. Rainfall amounts of 4-8 inches have been already reported in parts of Alabama, Mississippi and Louisiana. …On Sept. 4, Ike strengthened into a major hurricane, Category 4 on the Saffir-Simpson Scale, with maximum sustained winds near 145 mph. Ike is forecast to head west and may also affect the Bahamas. Behind Ike, on Sept. 2, the tenth tropical depression in the Atlantic Ocean basin was born. Tropical Depression 10 formed west of the African coast, so it has a long way to go before it has any impact on the U.S. or the Caribbean. By the late morning, that tropical cyclone became Tropical Storm Josephine. In August, Fay’s ten-day romp from the U.S. Southeast northward up the Appalachian Mountains seemed like a harbinger for September’s storms. Fay took her time going northward and dumped tremendous amounts of rain along the way. Melbourne Beach, Florida, received as much as 25.28 inches of rain. Other cities in various states reported high totals: Thomasville, Ga., reported 17.43 inches; Camden, Ala., received 6.85 inches; Beaufort, S.C., received 6.11 inches; Carthage, Tenn., reported 5.30 inches, and Charlotte, N.C., reported 5.90 inches. Fay was a perfect example of how weaker tropical storms can cause flooding inland … |
| 535 | …The topic has local interest because it has recently been observed that significant erosion is occurring on New Orleans area levees primarily caused by seepage-driven flow. Recently work began by the U.S. Army Corps of Engineers to plug in the shipping channel known as the Mississippi River-Gulf Outlet Canal (MRGO) that extends for 76 miles from New Orleans through wetlands to the Gulf of Mexico. The erosion of thousands of acres of cypress wetlands and marsh caused by the channel has been blamed for increased storm surge during Hurricane Katrina in 2005. Our theory would suggest that seepage caused by underwater flow will continue to erode and weaken the levee system around New Orleans, but the rate of this erosion should gradually slow with time, says Straub. Hopefully this research will aid the U.S. Army Corps of Engineers in identifying levees that need repair and assessing the lifespan of structures like the MRGO that are not planned for upkeep. Using fieldwork conducted in the Florida Panhandle, Straub and his fellow researchers were able to better understand the process of seepage erosion, which occurs when the re-emergence of groundwater at the surface shapes the Earth’s topography. The paper suggests that the velocity at which channel heads advance is proportional to the flux of groundwater to the heads. The researchers used field observations and numerical modeling to come up with the theory. … |
Appendix H AI Assistants Usage
We achieve most work by ourselves while using ChatGPT to partially provide coding help (e.g., grammar checks, package searches).

### v2

@@ 5.4. Ablation Study
5.5. Parameter Study
Hypercube-RAG combines sparse exact match and semantic retrieval strategies, as introduced in Section 3.3. We investigate the effect of the semantic retrieval strategy on the quality of LLM responses under varying similarity thresholds, denoted by . Figure 7 shows that performance improves with increasing threshold values when , suggesting that higher similarity scores are more effective for retrieving semantically relevant documents. However, when , the behavior of semantic RAG begins to gradually degrade to the sparse exact matching strategy, resulting in a decline in performance. This is primarily due to overly strict matching criteria, which reduce the number of retrieved documents and thus limit the available context.
5.6. Case Study
An additional case study in Figure 5 presents that Sparse BM25, Semantic-RAG (e5), and HippoRAG 2 fail to generate the correct answer due to inaccurate document retrieval. These methods tend to prioritize returning documents that are semantically similar in the embedding space. For example, semantic embedding methods tend to retrieve those documents that largely mention above-normal, near-normal, and below-normal hurricane season. However, these common terminologies may overshadow more specific, less frequent, but critical information, such as the location “Atlantic” and the organization “Climate Prediction Center”. In contrast, our Hypercube-RAG returns the correct answer. It accurately identifies and retrieves documents containing key elements of the query, including “Atlantic”, “hurricane season”, and “Climate Prediction Center”, corresponding to the hypercube dimensions, LOCATON, THEME, and ORGANIZATION, respectively. This retrieval approach intuitively prioritizes specific contextual information over abstract, ambiguous references such as above-, near-, and below-normal alone.
6. Discussion and Conclusions
In this work, we start by analyzing the strengths and limitations of existing RAG methods for scientific question-answering, which tend to perform poorly in either accuracy, efficiency, or explainability of the retrieval process. To retain their advantages while addressing key challenges, we introduce Hypercube, a novel structure representing knowledge in a compact, multi-dimensional space. Cube cell(s) are associated with compact, fine-grained labels extracted from the documents using LLMs.
Built on the hypercube, Hypercube-RAG is proposed to innovate the knowledge retrieval based on those cube labels. The cube label-based retrieval integrates sparse lexical and dense semantic strategies, making it accurate, efficient, and inherently explainable. We conduct experiments on three datasets across different domains comprehensively, and report the results using four retrieval metrics and two QA metrics, demonstrating that our method consistently outperforms existing baselines in both retrieval and question-answering accuracy. Additionally, we reveal a higher retrieval efficiency of Hypercube-RAG than existing approaches, and verify the explainable retrieval process. Overall, we conclude that Hypercube-RAG achieves high accuracy and efficiency, while providing interpretable provenance for retrieval.
Acknowledgements.
This work is supported by the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE), which is funded by the National Science Foundation (NSF) under award number: 2118329. Any opinions, findings, conclusions, or recommendations expressed herein are those of the authors and do not necessarily represent the views of NSF.References
- (1)
- Asai et al. (2023) Akari Asai, Zeqiu Wu, Yizhong Wang, Avirup Sil, and Hannaneh Hajishirzi. 2023. Self-rag: Learning to retrieve, generate, and critique through self-reflection. In The Twelfth International Conference on Learning Representations.
- Augenstein et al. (2024) Isabelle Augenstein, Timothy Baldwin, Meeyoung Cha, Tanmoy Chakraborty, Giovanni Luca Ciampaglia, David Corney, Renee DiResta, Emilio Ferrara, Scott Hale, Alon Halevy, et al. 2024. Factuality challenges in the era of large language models and opportunities for fact-checking. Nature Machine Intelligence 6, 8 (2024), 852–863.
- Chandrasekaran and Mago (2021) Dhivya Chandrasekaran and Vijay Mago. 2021. Evolution of semantic similarity—a survey. Acm Computing Surveys (Csur) 54, 2 (2021), 1–37.
- Cui et al. (2025) Hejie Cui, Alyssa Unell, Bowen Chen, Jason Alan Fries, Emily Alsentzer, Sanmi Koyejo, and Nigam Shah. 2025. TIMER: Temporal Instruction Modeling and Evaluation for Longitudinal Clinical Records. In Will Synthetic Data Finally Solve the Data Access Problem? https://openreview.net/forum?id=uBCAtA6M73
- Devezas and Nunes (2021) José Devezas and Sérgio Nunes. 2021. A review of graph-based models for entity-oriented search. SN Computer Science 2, 6 (2021), 437.
- Edge et al. (2024) Darren Edge, Ha Trinh, Newman Cheng, Joshua Bradley, Alex Chao, Apurva Mody, Steven Truitt, Dasha Metropolitansky, Robert Osazuwa Ness, and Jonathan Larson. 2024. From local to global: A graph rag approach to query-focused summarization. arXiv preprint arXiv:2404.16130 (2024).
- Feng et al. (2015) Wei Feng, Chao Zhang, Wei Zhang, Jiawei Han, Jianyong Wang, Charu Aggarwal, and Jianbin Huang. 2015. STREAMCUBE: Hierarchical spatio-temporal hashtag clustering for event exploration over the Twitter stream. In 2015 IEEE 31st international conference on data engineering. IEEE, 1561–1572.
- Guo et al. (2024) Zirui Guo, Lianghao Xia, Yanhua Yu, Tu Ao, and Chao Huang. 2024. Lightrag: Simple and fast retrieval-augmented generation. arXiv preprint arXiv:2410.05779 (2024).
- Gutiérrez et al. (2024) Bernal Jiménez Gutiérrez, Yiheng Shu, Yu Gu, Michihiro Yasunaga, and Yu Su. 2024. Hipporag: Neurobiologically inspired long-term memory for large language models. In The Thirty-eighth Annual Conference on Neural Information Processing Systems.
- Gutiérrez et al. (2025) Bernal Jiménez Gutiérrez, Yiheng Shu, Weijian Qi, Sizhe Zhou, and Yu Su. 2025. From RAG to Memory: Non-Parametric Continual Learning for Large Language Models. arXiv preprint arXiv:2502.14802 (2025).
- Han et al. (2024) Haoyu Han, Yu Wang, Harry Shomer, Kai Guo, Jiayuan Ding, Yongjia Lei, Mahantesh Halappanavar, Ryan A Rossi, Subhabrata Mukherjee, Xianfeng Tang, et al. 2024. Retrieval-augmented generation with graphs (graphrag). arXiv preprint arXiv:2501.00309 (2024).
- Huang et al. (2025) Lei Huang, Weijiang Yu, Weitao Ma, Weihong Zhong, Zhangyin Feng, Haotian Wang, Qianglong Chen, Weihua Peng, Xiaocheng Feng, Bing Qin, et al. 2025. A survey on hallucination in large language models: Principles, taxonomy, challenges, and open questions. ACM Transactions on Information Systems 43, 2 (2025), 1–55.
- Izacard et al. (2021) Gautier Izacard, Mathilde Caron, Lucas Hosseini, Sebastian Riedel, Piotr Bojanowski, Armand Joulin, and Edouard Grave. 2021. Unsupervised dense information retrieval with contrastive learning. arXiv preprint arXiv:2112.09118 (2021).
- Izacard et al. (2022) Gautier Izacard, Mathilde Caron, Lucas Hosseini, Sebastian Riedel, Piotr Bojanowski, Armand Joulin, and Edouard Grave. 2022. Unsupervised Dense Information Retrieval with Contrastive Learning. Transactions on Machine Learning Research (2022). https://openreview.net/forum?id=jKN1pXi7b0
- Ji et al. (2019) Xiaonan Ji, Han-Wei Shen, Alan Ritter, Raghu Machiraju, and Po-Yin Yen. 2019. Visual exploration of neural document embedding in information retrieval: Semantics and feature selection. IEEE transactions on visualization and computer graphics 25, 6 (2019), 2181–2192.
- Jin et al. (2025) Bowen Jin, Jinsung Yoon, Jiawei Han, and Sercan O Arik. 2025. Long-Context LLMs Meet RAG: Overcoming Challenges for Long Inputs in RAG. In The Thirteenth International Conference on Learning Representations. https://openreview.net/forum?id=oU3tpaR8fm
- Kadhim (2019) Ammar Ismael Kadhim. 2019. Term weighting for feature extraction on Twitter: A comparison between BM25 and TF-IDF. In 2019 international conference on advanced science and engineering (ICOASE). IEEE, 124–128.
- Kang et al. (2024) SeongKu Kang, Shivam Agarwal, Bowen Jin, Dongha Lee, Hwanjo Yu, and Jiawei Han. 2024. Improving retrieval in theme-specific applications using a corpus topical taxonomy. In Proceedings of the ACM Web Conference 2024. 1497–1508.
- Karpukhin et al. (2020) Vladimir Karpukhin, Barlas Oguz, Sewon Min, Patrick SH Lewis, Ledell Wu, Sergey Edunov, Danqi Chen, and Wen-tau Yih. 2020. Dense Passage Retrieval for Open-Domain Question Answering.. In EMNLP (1). 6769–6781.
- Krishna et al. (2021) Kalpesh Krishna, Aurko Roy, and Mohit Iyyer. 2021. Hurdles to Progress in Long-form Question Answering. In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Kristina Toutanova, Anna Rumshisky, Luke Zettlemoyer, Dilek Hakkani-Tur, Iz Beltagy, Steven Bethard, Ryan Cotterell, Tanmoy Chakraborty, and Yichao Zhou (Eds.). Association for Computational Linguistics, Online, 4940–4957. doi:10.18653/v1/2021.naacl-main.393
- Lee et al. (2025) Chankyu Lee, Rajarshi Roy, Mengyao Xu, Jonathan Raiman, Mohammad Shoeybi, Bryan Catanzaro, and Wei Ping. 2025. NV-Embed: Improved Techniques for Training LLMs as Generalist Embedding Models. In The Thirteenth International Conference on Learning Representations. https://openreview.net/forum?id=lgsyLSsDRe
- Lee et al. (2021) Kenton Lee, Ming-Wei Chang, and Kristina Toutanova. 2021. Learning Dense Representations of Phrases at Scale. In ACL.
- Lin et al. (2023) Xi Victoria Lin, Xilun Chen, Mingda Chen, Weijia Shi, Maria Lomeli, Richard James, Pedro Rodriguez, Jacob Kahn, Gergely Szilvasy, Mike Lewis, et al. 2023. Ra-dit: Retrieval-augmented dual instruction tuning. In The Twelfth International Conference on Learning Representations.
- Liu et al. (2024) Nelson F Liu, Kevin Lin, John Hewitt, Ashwin Paranjape, Michele Bevilacqua, Fabio Petroni, and Percy Liang. 2024. Lost in the middle: How language models use long contexts. Transactions of the Association for Computational Linguistics 12 (2024), 157–173.
- Maillard et al. (2021) Jean Maillard, Barlas Oguz, and Wen-tau Yih. 2021. Multi-Faith Retrieval for Question Answering. In NAACL.
- Mishra and Mittal (2021) Prakamya Mishra and Rohan Mittal. 2021. Neuralnere: Neural named entity relationship extraction for end-to-end climate change knowledge graph construction. In Tackling climate change with machine learning workshop at ICML.
- Peng et al. (2024) Boci Peng, Yun Zhu, Yongchao Liu, Xiaohe Bo, Haizhou Shi, Chuntao Hong, Yan Zhang, and Siliang Tang. 2024. Graph retrieval-augmented generation: A survey. arXiv preprint arXiv:2408.08921 (2024).
- Pipitone and Alami (2024) Nicholas Pipitone and Ghita Houir Alami. 2024. Legalbench-rag: A benchmark for retrieval-augmented generation in the legal domain. arXiv preprint arXiv:2408.10343 (2024).
- Pujari et al. (2023) Tejaskumar Pujari, Anil Kumar Pakina, and Anshul Goel. 2023. Explainable AI and governance: Enhancing transparency and policy frameworks through retrieval-augmented generation (RAG). IOSR Journal of Computer Engineering (2023).
- Robertson and Walker (1994) Stephen E Robertson and Steve Walker. 1994. Some simple effective approximations to the 2-poisson model for probabilistic weighted retrieval. In SIGIR’94: Proceedings of the Seventeenth Annual International ACM-SIGIR Conference on Research and Development in Information Retrieval, organised by Dublin City University. Springer, 232–241.
- Tao et al. (2018) Fangbo Tao, Chao Zhang, Xiusi Chen, Meng Jiang, Tim Hanratty, Lance Kaplan, and Jiawei Han. 2018. Doc2cube: Allocating documents to text cube without labeled data. In 2018 IEEE International Conference on Data Mining (ICDM). IEEE, 1260–1265.
- Trotman et al. (2014) Andrew Trotman, Antti Puurula, and Blake Burgess. 2014. Improvements to BM25 and language models examined. In Proceedings of the 19th Australasian Document Computing Symposium. 58–65.
- Wadden et al. (2020) David Wadden, Shanchuan Lin, Kyle Lo, Lucy Lu Wang, Madeleine van Zuylen, Arman Cohan, and Hannaneh Hajishirzi. 2020. Fact or fiction: Verifying scientific claims. arXiv preprint arXiv:2004.14974 (2020).
- Wang et al. (2025) Gelan Wang, Yu Liu, Shukai Liu, Ling Zhang, and Liqun Yang. 2025. REMFLOW: RAG-enhanced multi-factor rainfall flooding warning in sponge airports via large language model. International Journal of Machine Learning and Cybernetics (2025), 1–21.
- Wang et al. (2024b) Liang Wang, Nan Yang, Xiaolong Huang, Linjun Yang, Rangan Majumder, and Furu Wei. 2024b. Multilingual e5 text embeddings: A technical report. arXiv preprint arXiv:2402.05672 (2024).
- Wang et al. (2024a) Shijie Wang, Jiani Huang, Zhikai Chen, Yu Song, Wenzhuo Tang, Haitao Mao, Wenqi Fan, Hui Liu, Xiaorui Liu, Dawei Yin, et al. 2024a. Graph machine learning in the era of large language models (llms). ACM Transactions on Intelligent Systems and Technology (2024).
- Wang et al. (2022) Yi Wang, Wenhao Yu, Yining Wang, Chenguang Li, Soroush Liu, and Hoifung Poon. 2022. Text Embeddings by Weakly-Supervised Contrastive Pre-training. arXiv preprint arXiv:2212.03533 (2022).
- Wang et al. (2023) Zhaonan Wang, Bowen Jin, Wei Hu, Minhao Jiang, Seungyeon Kang, Zhiyuan Li, Sizhe Zhou, Jiawei Han, and Shaowen Wang. 2023. Geospatial Knowledge Hypercube. In Proceedings of the 31st ACM International Conference on Advances in Geographic Information Systems. 1–4.
- Wellawatte et al. (2025) Geemi P Wellawatte, Huixuan Guo, Magdalena Lederbauer, Anna Borisova, Matthew Hart, Marta Brucka, and Philippe Schwaller. 2025. ChemLit-QA: a human evaluated dataset for chemistry RAG tasks. Machine Learning: Science and Technology 6, 2 (2025), 020601.
- Xiong et al. (2021) Lee Xiong, Chenyan Xiong, Ye Li, Kwok-Fung Tang, Jialin Liu, Paul N. Bennett, Junaid Ahmed, and Arnold Overwijk. 2021. Approximate Nearest Neighbor Negative Contrastive Learning for Dense Text Retrieval. In International Conference on Learning Representations. https://openreview.net/forum?id=zeFrfgyZln
- Zhang et al. (2025a) Qinggang Zhang, Shengyuan Chen, Yuanchen Bei, Zheng Yuan, Huachi Zhou, Zijin Hong, Junnan Dong, Hao Chen, Yi Chang, and Xiao Huang. 2025a. A Survey of Graph Retrieval-Augmented Generation for Customized Large Language Models. arXiv preprint arXiv:2501.13958 (2025).
- Zhang et al. (2025b) Yunyi Zhang, Ruozhen Yang, Xueqiang Xu, Rui Li, Jinfeng Xiao, Jiaming Shen, and Jiawei Han. 2025b. Teleclass: Taxonomy enrichment and llm-enhanced hierarchical text classification with minimal supervision. In Proceedings of the ACM on Web Conference 2025. 2032–2042.
Appendix
Appendix A Hypercube
A.1. Hypercube Design and Construction
Comprehensively understanding the corpus is critical for pre-defining hypercube dimensions, because these dimensions structurally represent the knowledge (i.e., documents) in the multi-dimensional space. We follow these steps to determine the hypercube dimensions from the corpus.
A.2. Hypercube Structure (Dimensions)
Table 9 presents the hypercube dimensions used for different datasets.
| Datasets | Hypercube Dimensions |
| SciFact | “date”, ‘location”, “person”, “quantity”, “organizations_research_initiatives”, “medicine_health”, “genetics_biology”, “immunology_neuroscience”, “pharmacology” |
| LegalBench | “date”, “organization”, “quantity”, “location”, “person”, “company”, “money_finance”, “relationship”, “law_agreement_regulation”, “agreement_information” |
| SciDCC (Hurricane) | “location”, “person”, “event”, “organization”, “theme”, “date” |
| SciDCC (Geography) | “location”, “person”, “event”, “organization”, “theme”, “date” |
| SciDCC (Aging_Dam) | “dam”, “date”, “facility”, “location”, “organization”, “person”, “event”, “quant”, ’watershed’ |
Appendix B Prompt Templates
1. Comprehend the given question: understand what the question asks, how to answer it step by step, and all concepts, aspects, or directions that are relevant to each step.
2. Compose queries to retrieve documents for answering the question: each document is indexed by the entities or phrases occurred inside and those entities or phrases lie within the following dimensions: {dimensions}.
For each of the above dimensions, synthesize queries that are informative, self-complete, and mostly likely to retrieve target documents for answering the question. Note that each of your queries should be an entity or a short phrase and its associated dimension.
Example Input: “How much rainfall did Melbourne Beach, Florida receive from Tropical Storm Fay?”
Example Output:
1. query_dimension: ‘location’; query_content: ‘Atlantic’;
2. query_dimension: ‘event’; query_content: ‘Tropical Storm Fay’;
3. query_dimension: ’theme’; query_content: ‘rainfall’
Question: {question},
Retrieved documents: {retrieved document},
Output requirements:
- If the query asks the quantitative analysis, such as starting with ”How many”, ”How much”, ”How much greater”, ”How wide”, ”What percentage”, ”How far”, ”How long”, ”How old”, ”What portion”, ”What depth”, please directly output the quantitative answer as short as possible.
- If the query asks specific information, such as ”what percentage”, ”what likelihood”, ”which years”, ”who”, please directly output the answer without explanation or other information.
Question: {question},
Gold Answer: {gold_answer},
Predicted Answer: {predicted_answer}.
Evaluate criteria:
1. Correctness (0 or 1):
- Score 1 if the predicted answer is generally accurate and aligns with the key points in the reference answer.
- Score 0 if it has factual errors or misrepresents key information.
2. Completeness (0 or 1):
- Score 1 if the predicted answer covers the main points present in the reference answer.
- Score 0 if it misses essential information or fails to address the core of the question.
Output your evaluation in the following JSON format:
{“correctness”: int, “completeness”: int, “explanation”: your brief explanation}
Appendix C Discussion
In the following, we discuss different scenarios of hypercube usage: simple query, long query with multiple themes, and diverse query with shifted themes.
C.1. Case 1: simple query
How much rainfall did Melbourne Beach, Florida receive from Tropical Storm Fay? The relevant documents can be retrieved from one cube cell in a hypercube ¡Melbourne Beach, rainfall, Tropical Storm Fay ¿.
C.2. Case 2: long query with multiple themes
In cases where a query is very diverse, they may need to access multiple cube cells such that the query information can be covered as much as possible. For example, the diverse query could be “What consequences were caused by Tropical Storm Fay, such as how much rainfall Melbourne Beach and Palm Bay in Florida received, and how much monetary losses were caused?”. Those cube cells needed could be ¡Melbourne Beach, rainfall, Tropical Storm Fay ¿, ¡Palm Beach, rainfall, Tropical Storm Fay ¿, ¡Melbourne Beach, monetary losses, Tropical Storm Fay ¿, ¡Palm Beach, monetary losses, Tropical Storm Fay¿.
C.3. Case 3: diverse query with shifted themes
Given a more complicated query including shifted topics, e.g., “How did the monetary loss caused by Tropical Storm Fay impact the industrial layoff in Miami Downtown?”, cube cells across two hypercubes are needed ¡Miami Downtown, Industrial Layoff, Tropical Storm Fay ¿ and ¡Miami Downtown, Monetary loss, Tropical Storm Fay ¿ to get more relevant documents. These cross-theme queries are usually complex and need more dedicated hypercube design, which is considered for future exploration.